OpenStudy (anonymous):

Correct me if I'm wrong: 1. You have recently found a location for your bakery and have begun implementing the first phases of your business plan. Your budget consists of a $90,000 loan from your family and a $37,750 small business loan. These loans must be repaid in full within 10 years. a. What integer would represent your total budget? 127,750

OpenStudy (anonymous):

or should it be x=127,750?

OpenStudy (lilg132):

i think $127,750 should be a suitable answer

OpenStudy (anonymous):

ok, so i got that one right...so now the next part to this problem..let me know if need correction... b. You will use 20% of your budget to rent business space and pay for utilities. Write an algebraic expression that indicates how much money will be spent on business space and utilities. Do not solve. X=.20(127,750) 127,750*.20 multiply c. How much money will rent and utilities cost? Explain how you arrived at this answer. It will cost 25,550 total for rent and utilities. The answer was determined by multiplying the equation above. 127,750*.20.

OpenStudy (lilg132):

b is correct

OpenStudy (lilg132):

c is correct well done :)

OpenStudy (anonymous):

d. Imagine an investor has increased your budget by $26,000. The investor does not need to be repaid. Rather, he becomes part owner of your business. Will the investor contribute enough money to meet the cost of rent and utilities? Support your answer and write an equation or inequality that illustrates your answer. Yes. 26,000>25,550 e. This equation illustrates your remaining funds after paying for rent and utilities. How much money is left? Explain how you arrived at your answer. $37,750 + $90,000+ $26,000-0.20($90,000 + $37,750) = $450?

OpenStudy (anonymous):

not sure about E!

OpenStudy (lilg132):

do you have to write your own equation or did they give you the equation for question e) ?

OpenStudy (anonymous):

no, i didn't write that number for E. the only thing i had to come up with is the answer to what is leftover...so I came up with 450 dollars leftover

OpenStudy (anonymous):

E ties into all the above problems..

OpenStudy (anonymous):

and needing to explain how i got my answer to the leftover money is left.......

OpenStudy (lilg132):

well i got $128,200 left over

OpenStudy (anonymous):

hmm...can you tell me what you did to come up with that answer?

OpenStudy (lilg132):

so what you do is -0.20 multiplied by 90,000 which is -18,000 and -0.20 multiplied by 37,750 which is -7550 so you do -18000 - 7550 = -25550 and then you do -25,550 + 26,000 + 90,000 + 37,750 = $128,200

OpenStudy (anonymous):

aaah,...i was confused on that one, wasn't sure..

OpenStudy (anonymous):

going to post another to make sure its right..new post!!